# Dissipation-dispersion analysis of fully-discrete implicit discontinuous Galerkin methods and application to stiff hyperbolic problems

### v2

@@ 5.1.1 Convergence test
$0.1$ $20$ $21.47e{-}$ 2 $2.80$ $22.17e{-}$ 2 $2.75$ $0.2$ $1.38$ $0.43$ $1.97$ $1.27$ $1.68$
$5.00\cdot 10^{-2}$ $40$ $32.11e{-}$ 3 $2.80$ $33.67e{-}$ 3 $2.57$ $23.48e{-}$ 2 $2.49$ $0.12$ $1.84$ $0.45$ $1.51$
$2.50\cdot 10^{-2}$ $80$ $43.99e{-}$ 4 $2.40$ $46.29e{-}$ 4 $2.54$ $35.96e{-}$ 3 $2.55$ $22.32e{-}$ 2 $2.38$ $0.12$ $1.91$
$1.25\cdot 10^{-2}$ $160$ $55.98e{-}$ 5 $2.74$ $41.25e{-}$ 4 $2.33$ $47.81e{-}$ 4 $2.93$ $33.23e{-}$ 3 $2.84$ $22.28e{-}$ 2 $2.39$
$6.25\cdot 10^{-3}$ $320$ $67.24e{-}$ 6 $3.05$ $54.39e{-}$ 5 $1.51$ $59.87e{-}$ 5 $2.98$ $44.00e{-}$ 4 $3.01$ $33.15e{-}$ 3 $2.85$
$3.13\cdot 10^{-3}$ $640$ $63.64e{-}$ 6 $3.59$
$1.56\cdot 10^{-3}$ $1{,}280$ $74.63e{-}$ 7 $2.97$
@@ 5.1.3 Burgers’ equation
on the periodic domain $x\in[0,2]$, and up to the final time $t=2$, i.e. after shock formation. Then, we test the Burgers’ equation on the discontinuous initial condition (5.3), on the periodic domain $x\in[-1,1]$, and up to the final time $t=0.5$. Figures 10 and Figure 11 present the results of different DIRK-DG schemes. All solutions are computed with a tolerance of $10^{-5}$ for the GMRES solver, and with a maximum number of iterations limited to $100$. On average, the GMRES solver converges within 10 iterations for the second-order schemes and approximately 15 iterations for the third-order schemes, with slightly higher values at initial time steps. A detailed technical study of the iteration behavior is beyond the scope of this work.
@@ 5.2.1 Convergence test
In order to test the accuracy of the implicit scheme, we consider $\kappa=0$ and $\kappa=4$ and run the simulation with the following ratio between the explicit and the implicit time-step
@@
Table 13: Experimental orders of convergence computed on the Euler system (5.7) with initial condition (5.8).
@@
$N$ $L^{1}$ error rate $L^{1}$ error rate
$40$ $22.83e{-}$ 2
$80$ $21.15e{-}$ 2 $1.30$
$160$ $32.86e{-}$ 3 $2.01$ $28.37e{-}$ 2
$320$ $47.19e{-}$ 4 $1.99$ $22.38e{-}$ 2 $1.81$
$640$ $41.92e{-}$ 4 $1.91$ $36.16e{-}$ 3 $1.95$
$1{,}280$ $32.03e{-}$ 3 $1.60$
@@
$N$ $L^{1}$ error rate $L^{1}$ error rate
$40$ $21.69e{-}$ 2
$80$ $33.41e{-}$ 3 $2.31$
$160$ $45.17e{-}$ 4 $2.72$ $21.07e{-}$ 2
$320$ $56.88e{-}$ 5 $2.91$ $33.46e{-}$ 3 $1.63$
$640$ $68.39e{-}$ 6 $3.04$ $44.54e{-}$ 4 $2.93$
$1{,}280$ $55.28e{-}$ 5 $3.10$
Table 13 presents the experimental orders of convergence computed for the Euler system under two different schemes: the second-order scheme ${\mathsf{DG}}_{2}\mathsf{DIRK}_{22}^{\mathsf{Lstab}}$ in 13(a) and the third-order scheme ${\mathsf{DG}}_{3}\mathsf{DIRK}_{33}^{\mathsf{Lstab}}$ in Table 13(b). Both tables provide data for different levels of grid refinement and for two different values of $\kappa$, which defines the level of stiffness.
@@ 5.2.2 Stiff Riemann problems
(a) Second-order schemes. Left: density solution of the non-symmetric expansion problem. Right: density solution of the colliding flows problem.
(b) Third-order schemes. Left: density solution of the non-symmetric expansion problem. Right: density solution of the colliding flows problem.
@@ 5.2.3 Comparison with the third-order Quinpi finite-volume scheme
(a) Non-symmetric expansion problem. The left-most panel shows the density solution. The other panels show zooms on the $1$-rarefaction, the contact and $2$-rarefaction wave, respectively.
(b) Colliding flow problem. The left-most panel shows the density solution. The other panels show zooms on the $1$-shock the contact and $2$-shock wave, respectively
Figure 16: Comparison between the third-order DIRK-DG implicit scheme and the third-order finite volume implicit scheme from [44]. For the non-symmetric expansion problem, N=266N=266 cells are used for the ${\mathsf{DG}}_{3}\mathsf{DIRK}_{33}^{\mathsf{Lstab}}$ scheme and $N=800$ cells for Quinpi. For the colliding flow problem, $N=666$ cells are used for the ${\mathsf{DG}}_{3}\mathsf{DIRK}_{33}^{\mathsf{Lstab}}$ scheme and $N=2000$ cells for Quinpi. The number of cells is chosen to ensure that both schemes are computed with the same degrees of freedom.
Figure 16 provides a detailed comparison between the third-order DIRK-DG implicit scheme and the third-order finite volume implicit scheme from [44] on the non-symmetric expansion problem (Figure 16(a)) and the colliding flow problem (Figure 16(b)). For the non-symmetric expansion, the simulations were carried out using $N=266$ cells for the ${\mathsf{DG}}_{3}\mathsf{DIRK}_{33}^{\mathsf{Lstab}}$ scheme and $N=800$ cells for Quinpi. In the colliding flow scenario, $N=666$ cells were used for the ${\mathsf{DG}}_{3}\mathsf{DIRK}_{33}^{\mathsf{Lstab}}$ scheme and $N=2000$ cells for Quinpi. The choice of cell numbers was made to ensure that both schemes operated with the same degrees of freedom, facilitating a fair comparison between them.
@@ 6 Conclusions
Our results highlight the strengths of the newly developed implicit discontinuous Galerkin schemes. Specifically, these methods demonstrate the capability to accurately capture slow material waves in stiff problems, effectively overcoming the stringent stability constraints, based on fast acoustic waves, that typically limit explicit schemes. This advantage is particularly evident in scenarios where the phenomena of interest propagate at speeds much slower than the fastest wave speeds, allowing the implicit schemes to operate with time-steps much larger than those required by explicit methods, while still achieving better accuracy on the phenomenon of interest.
@@ Acknowledgments
This work was carried out within the Ateneo Sapienza project 2023 “Modeling, numerical treatment of hyperbolic equations and optimal control problems”.
GV acknowledge support from PNRR-MUR project “Italian Research Center on High Performance Computing, Big Data and Quantum Computing”.
@@
No conflict of interest is extant in the present work.

### v3

@@ 5.1.1 Convergence test
$0.1$ 2020 1.47e21.47e{-}2 $2.80$ 2.17e22.17e{-}2 $2.75$ $0.2$ $1.38$ $0.43$ $1.97$ $1.27$ $1.68$
$5.00\cdot 10^{-2}$ $40$ 2.11e32.11e{-}3 $2.80$ 3.67e33.67e{-}3 $2.57$ 3.48e23.48e{-}2 $2.49$ $0.12$ $1.84$ $0.45$ $1.51$
$2.50\cdot 10^{-2}$ $80$ 3.99e43.99e{-}4 $2.40$ 6.29e46.29e{-}4 $2.54$ 5.96e35.96e{-}3 $2.55$ 2.32e22.32e{-}2 $2.38$ $0.12$ $1.91$
$1.25\cdot 10^{-2}$ $160$ 5.98e55.98e{-}5 $2.74$ 1.25e41.25e{-}4 $2.33$ 7.81e47.81e{-}4 $2.93$ 3.23e33.23e{-}3 $2.84$ 2.28e22.28e{-}2 $2.39$
$6.25\cdot 10^{-3}$ $320$ 7.24e67.24e{-}6 $3.05$ 4.39e54.39e{-}5 $1.51$ 9.87e59.87e{-}5 $2.98$ 4.00e44.00e{-}4 $3.01$ 3.15e33.15e{-}3 $2.85$
$3.13\cdot 10^{-3}$ $640$ 3.64e63.64e{-}6 $3.59$
$1.56\cdot 10^{-3}$ 1,2801{,}280 4.63e74.63e{-}7 $2.97$
@@ 5.1.3 Burgers’ equation
on the periodic domain $x\in[0,2]$, and up to the final time $t=2$, i.e. after shock formation. Then, we test the Burgers’ equation on the discontinuous initial condition (5.3), on the periodic domain $x\in[-1,1]$, and up to the final time $t=0.5$. Figures 10 and Figure 11 present the results of different DIRK-DG schemes. In these numerical tests $\delta$ is chosen as $\delta\approx r/(2p+1)$. All solutions are computed with a tolerance of $10^{-5}$ for the GMRES solver, and with a maximum number of iterations limited to $100$. On average, the GMRES solver converges within 10 iterations for the second-order schemes and approximately 15 iterations for the third-order schemes, with slightly higher values at initial time steps. A detailed technical study of the iteration behavior is beyond the scope of this work.
@@ 5.2.1 Convergence test
In order to test the accuracy of the implicit scheme, we consider $\kappa=0$ and $\kappa=2$ and run the simulation with the following ratio between the explicit and the implicit time-step
@@
These choices imply the following implicit time-steps
$\Delta t_{\text{imp}}=\begin{cases}3.563,&\kappa=0,\ p=1\\ 2.138,&\kappa=0,\ p=2\\ 2.364,&\kappa=2,\ p=1\\ 1.419,&\kappa=2,\ p=2.\end{cases}$
@@
Table 13: Experimental orders of convergence computed on the Euler system (5.7) with initial condition (5.8). The errors are computed on the density.
@@
NN $L^{1}$ error rate $L^{1}$ error rate
$40$ 2.83e22.83e{-}2
$80$ 1.15e21.15e{-}2 $1.30$
$160$ 2.86e32.86e{-}3 $2.01$ 8.37e28.37e{-}2
$320$ 7.19e47.19e{-}4 $1.99$ 2.38e22.38e{-}2 $1.81$
$640$ 1.92e41.92e{-}4 $1.91$ 6.16e36.16e{-}3 $1.95$
1,2801{,}280 2.03e32.03e{-}3 $1.60$
@@
NN $L^{1}$ error rate $L^{1}$ error rate
$40$ 1.69e21.69e{-}2
$80$ 3.41e33.41e{-}3 $2.31$
$160$ 5.17e45.17e{-}4 $2.72$ 1.07e21.07e{-}2
$320$ 6.88e56.88e{-}5 $2.91$ 3.46e33.46e{-}3 $1.63$
$640$ 8.39e68.39e{-}6 $3.04$ 4.54e44.54e{-}4 $2.93$
1,2801{,}280 5.28e55.28e{-}5 $3.10$
Table 13 presents the experimental orders of convergence computed for the Euler system under two different schemes: the second-order scheme ${\mathsf{DG}}_{2}\mathsf{DIRK}_{22}^{\mathsf{Lstab}}$ in 13(a) and the third-order scheme ${\mathsf{DG}}_{3}\mathsf{DIRK}_{33}^{\mathsf{Lstab}}$ in Table 13(b). Both tables provide data for different levels of grid refinement and for two different values of $\kappa$, which defines the level of stiffness. The errors are computed on the density profile.
@@ 5.2.2 Stiff Riemann problems
(a) Second-order schemes. Left: density solution in the $x$-$t$ diagram of the non-symmetric expansion problem. Right: density solution in the $x$-$t$ diagram of the colliding flows problem. The black lines represent the wave structures.
(b) Third-order schemes. Left: density solution in the $x$-$t$ diagram of the non-symmetric expansion problem. Right: density solution in the $x$-$t$ diagram of the colliding flows problem. The black lines represent the wave structures.
@@ 5.2.3 Comparison with the third-order Quinpi finite-volume scheme
(a) Non-symmetric expansion problem. The left-most panel shows the density solution. The other panels show zooms on the 11-rarefaction, the contact and 22-rarefaction wave, respectively.
(b) Colliding flow problem. The left-most panel shows the density solution. The other panels show zooms on the 11-shock the contact and 22-shock wave, respectively
Figure 16: Comparison between the third-order DIRK-DG implicit scheme and the third-order finite volume implicit scheme from [44]. For the non-symmetric expansion problem, N=266 cells are used for the ${\mathsf{DG}}_{3}\mathsf{DIRK}_{33}^{\mathsf{Lstab}}$ scheme and $N=800$ cells for Quinpi. For the colliding flow problem, $N=666$ cells are used for the ${\mathsf{DG}}_{3}\mathsf{DIRK}_{33}^{\mathsf{Lstab}}$ scheme and $N=2000$ cells for Quinpi. The number of cells is chosen to ensure that both schemes are computed with comparable number of degrees of freedom.
Figure 16 provides a detailed comparison between the third-order DIRK-DG implicit scheme and the third-order finite volume implicit scheme from [44] on the non-symmetric expansion problem (Figure 16(a)) and the colliding flow problem (Figure 16(b)). For the non-symmetric expansion, the simulations were carried out using $N=266$ cells for the ${\mathsf{DG}}_{3}\mathsf{DIRK}_{33}^{\mathsf{Lstab}}$ scheme and $N=800$ cells for Quinpi. In the colliding flow scenario, $N=666$ cells were used for the ${\mathsf{DG}}_{3}\mathsf{DIRK}_{33}^{\mathsf{Lstab}}$ scheme and $N=2000$ cells for Quinpi. The choice of cell numbers was made to ensure that both schemes operated with comparable number of degrees of freedom, facilitating a fair comparison between them.
@@ 6 Conclusions
Our results highlight the strengths of the newly developed implicit discontinuous Galerkin schemes. Specifically, these methods demonstrate the capability to accurately capture slow material waves in stiff problems, effectively overcoming the stringent stability constraints, based on fast acoustic waves, that typically limit explicit schemes. This advantage is particularly evident in scenarios where the phenomena of interest propagate at speeds much slower than the fastest wave speeds, allowing the implicit schemes to operate with time-steps much larger than those required by explicit methods, while still achieving better accuracy on the phenomenon of interest. At the same time, our results confirm that this advantage comes at the cost of reduced accuracy for fast acoustic waves when large time steps are employed, a limitation that is intrinsic to implicit time integration. Such a trade-off is acceptable in applications where the phenomena of interest are dominated by slow dynamics.
@@
This work was carried out within the Ateneo Sapienza Project 2023 “Modeling, numerical treatment of hyperbolic equations and optimal control problems” funded by Sapienza University of Rome.
GV acknowledge support from PNRR-MUR (Ministry of University and Research) project “Italian Research Center on High Performance Computing, Big Data and Quantum Computing”.
@@
Conflict of interest
There is no conflict of interest in the present work.
Data Availability Statement
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.